OpenStudy (mitchi35):

So when trying to graph a circle, and say for instance you have...Graph the circle (x-4)2 (square root) +(y-2)2=25. Because they both are negatives, I suppose. Does that mean that x-coordinate of the center would be positive 4? And the y-coordinate of the center would be positive 2? And would that be the other way around if they were both negatives? Or one negative one positive? Will the coordinates always be the opposite?

Nnesha (nnesha):

yes right if it says (x-3)^2 then x-coordinate would be at positive 3 opposite sign general fomr equation \[\large\rm (x-h)^2+(y-k)^2=r^2\] \[\huge\rm (x-3)^2+(y+4)^2=9\] is same as \[(x-(+3))^2 +(y-(-4))^2=9\] center would be at (3,-4)

Nnesha (nnesha):

`Graph the circle (x-4)2 (square root) +(y-2)2=25` square root ???

OpenStudy (mitchi35):

Sorry, yeah I meant ^2. But thank you very much!

Nnesha (nnesha):

np :=)good work! :=) welcome to openstudy! btw

OpenStudy (mitchi35):

Thanks! :)